A 12.5855 g sample of an unknown compound (X) is dissolved in 32.0000 g of water. The resulting solution has a freezing point of – 4.3oC. What is the molar mass of X?   Kf = 1.858 oC/m.
Answer #1

ato a ix K4 x molality A7s + de pression in freezing point... in point of solvent- freezing point of solution =DI. ist freeziAt It (solvent) - If (solution ) 044 = 0°C - (-4-3%) ats = 4.3°C A Tt = Ksx mollity 4.3% = 1.858°/mx molality molality=2-3143